OpenStudy (anonymous):

If 7a + 6g = 31.40 and 5a + 10g = 37.00 what do variables a and g equal?

OpenStudy (hba):

There are a number of ways to solve this, what we could do is get a from equation 1 and subsitute it in equation 2 so we'll find g then we can put that value to find a

OpenStudy (anonymous):

Ok that sounds good can you show me how?

OpenStudy (anonymous):

from 5a+10g=37, rearrange to 5a=37/10g divide both sides by 5, a=37/50g just plug that a into the other equation and solve for g then once you have g plug that in to either one and solve for a

OpenStudy (hba):

lIke @14mdaz explained